how many total atoms are in .220g of P2O5?
1
Expert's answer
2019-09-24T05:54:22-0400

At the first let's find the molar mass of "{P_2}{O_5}" - it can be calculated as the sum of atomic weights taking into account the number of atoms of each type in the molecule

"{M_{{P_2}{O_5}}} = 2{M_P} + 5{M_O} \approx 2 \cdot 16[\frac{{\rm{g}}}{{{\rm{mol}}}}] + 5 \cdot 30.97[\frac{{\rm{g}}}{{{\rm{mol}}}}] = 141.94[\frac{{\rm{g}}}{{{\rm{mol}}}}]"

(atomic weights of atoms "P" and "O" can be found in periodic table and respectively equal to "30.97[\frac{{\rm{g}}}{{{\rm{mol}}}}]" and "16[\frac{{\rm{g}}}{{{\rm{mol}}}}]")

The number of moles in "0.220[{\rm{g}}]" of "{P_2}{O_5}" can be found as


"\nu = \frac{m}{{{M_{{P_2}{O_5}}}}} = \frac{{0.220[{\rm{g}}]}}{{141.94[\frac{{\rm{g}}}{{{\rm{mol}}}}]}} \approx 1.55 \cdot {10^{ - 3}}[{\rm{mol}}]"

Now let's use that there are "{N_A} \approx 6.022 \cdot {10^{23}}[\frac{1}{{{\rm{mol}}}}]" (particles per 1 mole). In our case it will be the number of molecules. But to find the total number of atoms we can notice that there are "7" atoms in "1" molecule of "{P_2}{O_5}" (2 atoms of "P" and 5 atoms of "O"). Thus the total number of atoms is


"N = 7\nu {N_A} = 7 \cdot 1.55 \cdot {10^{ - 3}}[{\rm{mol}}] \cdot 6.022 \cdot {10^{23}}[\frac{1}{{{\rm{mol}}}}] \approx 6.53 \cdot {10^{21}}"